A local theatre sold 500 seats for an upcoming play. Ticket prices for adults are $62.00 and ticket prices for children are $32.00. If the theatre made $26,140.00 for the show, how many adult and children attended?
Answer by stanbon(75887)

A local theatre sold 500 seats for an upcoming play. Ticket prices for adults are $62.00 and ticket prices for children are $32.00. If the theatre made $26,140.00 for the show, how many adult and children attended?
-----
Quantity:: a + c = 500 tickets
Value:: 62a + 32c = 26140
----------
Modify for elimination::
31a + 31c = 31*500
31a + 16c = 13070
------
Subtract and solve for "c":
15c = 2430
c = 162 (# of child tickets sold)
----------
a = 500-162 = 338 (# of adult tickets sold)
---------------
Cheers,
Stan H.
----------